# Quantum thermocouples: nonlocal conversion and control of heat in semiconductor nanostructures

### v1

@@ I Introduction
The thermocouple is the minimal thermoelectric device able to use the Seebeck effect [1] to transduce a heat flow into an electric response measurable in an isothermal conductor. It is useful for the autonomous conversion of environmental (or waste) heat into useful power [2, 3], or as temperature sensors when coupled to a voltmeter. The basic principle of a macroscopic thermocouple is sketched in Fig. 1(a). Two junctions separate the conductor, at a base temperature $T_{C}$, from a region that is in contact with a heat source at temperature $T_{H}$. If the two junctions have opposite thermoelectric response (e.g., one being n-type and the other p-type), the electron-hole excitations, thermalized at a temperature $T_{M}$ in the central region, are spatially separated to opposite contacts. This can then either generate a current or be measured as a voltage difference which depends on the difference of the linear Seebeck coefficient of each junction: $S_{p}-S_{n}$. The reverse process is also possible: driving a current through the conductor can be used to extract heat from the central region, in which case one has a Peltier refrigerator.
@@
Reducing the size of the thermocouple to nanometer scales [4] comes with a number of key differences. The first to be noted was that the reduced dimensionality provides the required spectral properties for efficient thermoelectric response [5, 6], which can also be easily tuned with gate voltages, as illustrated in Fig. 1(b) for a resonant-tunneling barrier. At this scale, it becomes difficult to maintain large temperature differences. However, one benefits from the system response easily becoming nonlinear. From the point of view of the physical processes, the size of the region in contact with the heat source can be of the order of or smaller than the carrier thermalization length. This has two possible consequences: the thermoelectric response may become nonlocal [7, 8] (not determined by the series response of the junctions) or determined by nonthermal states [9], as represented in Fig. 1(c). Finally, at low enough temperatures quantum properties arising from electrons propagating as waves in conductors smaller than the phase coherence length start to contribute, leading to interference effects [10, 11]. Also, single particle interactions (electron-electron or electron-photon) can be relevant [12]. This is the regime where we focus our interest in this review. Properties from quantum materials (e.g., superconductors or edge states in quantum Hall regimes) can also be exploited [8].
@@
While the electronic transport through nanostructures has been intensively investigated during the last decades both experimental [13] and theoretically [14], progress in the control of thermal and thermoelectric flows was somehow delayed, mainly due to the difficulty in controlling and measuring minute temperature differences in an experiment and eliminating phonon drag contributions. This has however changed in the last years [15, 16, 17]. The interest is not restricted to the possibility to generate useful thermodynamic outputs but also to use the thermoelectric response as sensitive probes for transport mechanisms, in particular involving the difficult to measure energy flow. Controlled thermoelectric measurements have been achieved in 2DEGs [18, 19], quantum point contacts [20, 21, 22, 23], quantum dots [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], single-electron transistors [37], nanowires [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], scanning probes [49, 50], molecular junctions [51, 52, 53, 54], superconducting tunnel junctions [55, 56, 57] or magnetic wires [58], also in cold atom gases [59]. In particular, three-terminal quantum dot arrangements in thermocouple geometries have been realized using electron-electron interactions [60, 61, 62], hot cavities [63], electron-phonon [64, 65] or electron-photon interactions [66, 67] to couple the conductor to the heat source. Multiterminal quantum Hall edge states [68, 69], (local) Joule heating in graphene [70] and normal-superconductor interfaces [71] have also been considered. Interestingly, similar responses have been measured when coupling the conductor to nonequilibrium fluctuations [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85], which can be related to energy quanta emitted by relaxing electrons [86].
@@
This progress has been accompanied by theoretical works proposing the multiterminal thermoelectric response to take advantage of the separation of thermal and particle currents [87, 88, 89] or to access properties not available in the electrical conductance, for instance nonabelian statistics of carriers [90], odd-frequency superconducting correlations [91], crossed-Andreev reflection [92, 93, 94], correlations between edge states [95] or details of interfaces between superconductors and topological materials [96, 97, 98]. At the same time, questions related to quantum thermodynamics, e.g. the efficiency of quantum heat engines [99, 100], or the role of fluctuations [101, 102, 103, 104, 105, 106] and of information [87, 107, 108, 109, 110, 111, 112, 113] have found in the thermoelectric effect a suitable platform to be addressed.
@@
Particularly relevant to our discussion in this review is the configuration where a quantum system is coupled to two electronic terminals at the same temperature, with the third terminal injecting only heat. Under the appropriate broken symmetries, the injected heat is rectified into a particle current. Initial proposals considered electron-phonon coupling [114] and electron-electron interactions [87] as possible mechanisms for the transfer of heat (later realized experimentally [60, 61, 62, 64]), followed by configurations exploiting inelastic scattering [115] (realized with quantum dots [63]), magnons [116] or photons [117, 118, 119]. The variety of transport configurations considered is large, including capacitively coupled quantum dots [87, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129], molecules [114, 130, 131], nanowires [132], resonant tunneling [115, 133, 134, 135, 136, 137, 138, 139], edge states [140, 141, 142, 143, 144, 145, 146, 95, 147], beam splitters [148, 149, 104, 150, 151], quantum dot arrays [152, 153, 154, 155, 88, 156, 157, 158, 159, 160], and tunnel [117, 161, 162], p-n [163] or Josephson [164] junctions. Three terminal geometries have also been considered for exploring the thermoelectric efficiency bounds when one of the terminals is a probe [165, 166, 167, 168, 169, 170, 171, 172]. In the case that the resource consists on several terminals, nonthermal configurations [123] or thermal drag effects [173, 174, 175] can be exploited.
@@ II Multiterminal particle and heat transport
We are interested in configurations as the one depicted in Fig. 1(c), with multiple (more than two) terminals allowing for particle ($I_{i}$) and heat ($J_{i}$) currents, which will be positive when flowing out of their reservoirs, labeled by $i$ or $j$. The system is taken out of equilibrium by differences of the reservoir electrochemical potentials, $\mu_{i}$, and temperatures, $T_{i}$. Allowing the system to be composed of electrically isolated regions (i.e., in the case of having capacitively-coupled conductors), labeled by $\alpha$, particle and energy conservation are expressed as $\sum_{i}I_{\alpha_{i}}=0$ and $\sum_{i}J_{i}+\sum_{\alpha}P_{\alpha}=0$, where $P_{\alpha}=-\sum_{i}\mu_{i}I_{\alpha_{i}}$ is the dissipated power in conductor $\alpha$. The system will perform as a thermodynamic engine when power is generated ($P_{\alpha}>0$) in some conductor $\alpha$, as a refrigerator when heat flows out of a cold reservoir ($J_{i}>0$), or as a heat pump when heat flows into a hot reservoir ($J_{i}<0$). In a multiterminal setup, whether a terminal is considered “hot” or “cold” depends on its temperature being larger or smaller than a certain reference temperature $T_{0}$ [177]. Differently from macroscopic thermocouples though, fluctuating systems that are smaller than the electronic thermalization length may contain regions without a well defined temperature [123].
Our discussion is not restricted to the linear response regime, however considering this regime gives useful intuition. Linearizing the particle and heat currents with respect to variations of the electrochemical potentials and temperatures, we get the elements of the Onsager matrix:
@@
Multiterminal heat engines have obviously multiple possible resources and multiple possible thermodynamically useful outputs, that can even occur simultaneously [134, 182, 183, 145, 184, 177, 146, 185, 159, 186, 159, 187]. Hence, the usual definition of efficiency in terms of the ratio of output power and absorbed heat current needs to be revisited. Think for instance in the configuration of Fig. 1(c): uncareful definitions of the efficiency may diverge when power is generated by a current beween terminals $L$ and $R$, while the heat absorbed by the system, $J_{H}+J_{C}$, vanishes. To have a thermodynamically consistent definition of the engine performance, it is useful to consider the change in the nonequilibrium free energy of the reservoirs, $\dot{F}_{i}=-I_{i}\mu_{i}+J_{i}(T_{0}-T_{i})/T_{i}$ [177, 146], with the second law imposing that $\sum_{i}\dot{F}_{i}\leq 0$, i.e., free energy is consumed in dissipation. The efficiency is then written by selecting the positive electric and thermal contributions and dividing them by the negative ones [177]:
@@
where $\sum_{i}^{\pm}x_{i}=\sum(x_{i}\pm|x_{i}|)/2$. Generated power is positive, dissipated power is consumed therefore negative. Extracting/injecting heat from/into cold/hot reservoirs (i.e., cooling/pumping) are positive, damped heat is negative. As an example, consider the basic three-terminal thermocouple, with one hot reservoir injecting heat but no particles ($I_{H}=0$, $J_{H}>0$) and the other two terminals being at the same temperature ($T_{L}=T_{R}=T_{0}$) and supporting a particle current ($I_{L}+I_{R}=0$). Then if power is generated by the current flowing against the electrochemical potential difference, $I_{L}(\mu_{R}-\mu_{L})$, we recover $\eta=P/[J_{H}\eta_{C}]$, with the Carnot efficiency $\eta_{C}=1-T_{0}/T_{H}$.
@@ III Noninteracting electrons
A simple and elegant description of multiterminal quantum transport is given by scattering theory [189], as long as interactions admit a mean field approach [190]. Reservoirs at different electrochemical potentials $\mu_{i}$ and temperature $T_{i}$ inject electrons into the terminals of the conductor, which is treated as a wave scatterer. The scattering matrix ${\cal S}_{ji}(E)$ represents the amplitude for electrons injected with energy $E$ from terminal $i$ to be transmitted and absorbed by reservoir $j$. The success of the theory originates in the understanding of the properties and symmetries of the linear response regime [191, 192]. However, particle ($I_{i}$) and heat ($J_{i}$) currents can be more generally defined in terms of the occupation of the different terminals, given by the fermionic reservoir distribution function, $f_{i}(E)=1/\{1+\exp(E-\mu_{i})/k_{\text{B}}T_{i}]\}$:
[…]
@@
where $h$ is the Planck constant, the factor 2 is due to spin degeneracy and ${\cal T}_{ji}(E)=|{\cal S}_{ji}(E)|^{2}$ are the transmission probabilities. The unitarity of the scattering matrix preserves particle and energy conservation [8] as expressed in Sec. II. With this convention (particle and heat currents are positive when flowing out of reservoirs), the linear coefficients in Eq. (7) can be written as $G_{ij}=g_{ij}^{(0)}$, $L_{ij}=g_{ij}^{(1)}/T$, $M_{ij}=g_{ij}^{(1)}$, $K_{ij}=g_{ij}^{(2)}/T$, with [193]
[…]
@@
In the limit $\mu\gg k_{\text{B}}T$, these expressions show that the thermoelectric response requires that the scattering problem includes some explicit energy dependence i.e., that the nanostructure acts as a filter to break electron-hole symmetry. If ${\cal T}_{ij}=1$, one gets $G_{ij}=2/h$, $L_{ij}=M_{ij}=0$ and $K_{ij}=2q_{H}$, twice the quantum of thermal conductance, $q_{H}=\pi^{2}k_{\text{B}}^{2}T/3h$ [194]. The fulfillment of the Onsager reciprocity relations is hence related to the symmetries of the scattering matrix under time reversal: ${\cal T}_{ij}(B)={\cal T}_{ji}(-B)$ [192]. The coefficients plotted in Fig. 1(b) correspond to a Lorentzian transmission probability usually found in quantum dots [24, 25, 29].
[…]
@@
corresponding to the generated current and the probe condition, respectively. Note that the elastic transmission, ${\cal T}_{LR}$, does not contribute, and the generated zero bias current is entirely due to the inelastic scattering of electrons at the probe terminal. By inspection of Eq. (12), we see that for the system to work as a thermocouple able to generate a finite current, $I\neq 0$, it is required that the two transmissions are energy-dependent (in order to generate a thermoelectric response via the broken electron-hole symmetry), and that this dependence also breaks inversion (left-right) symmetry, i.e., $\mathcal{T}_{HL}(E)\neq c\mathcal{T}_{HR}(E)$, with $c$ a positive number. Otherwise, the two integrals become proportional, so the probe condition imposes $I=0$. This implies that the symmetry breaking must be a property of the conductor, not of the coupling to the heat source.
@@
To generate a finite power, the particle current must flow against a finite bias, $\mu_{L}\neq\mu_{R}$. In this case, the elastic transmissions ${\cal T}_{LR}(E)$ and ${\cal T}_{RL}(E)$ introduce contributions that are independent of the heat source (assuming that we can neglect nonlinear effects on the internal potential of the conductor [198, 199, 200, 201, 202]) and therefore are detrimental to the device performance [176]. Therefore power will be enhanced when the heat source is connected in series between the two conductor terminals [115], so all transported electrons are forced to thermalize at the probe. In analogy with the macroscopic thermocouple, the effect is enhanced if the two contributions to the current in Eq. (12) have opposite signs, i.e., one is dominated by electrons and the other by holes. This can be done by placing scatterers that separate the heat source from the two conductor terminals having sufficiently sharp spectral features antisymmetric with respect to the electrochemical potential, e.g., quantum dots [115]. It was found that in order to optimize the efficiency for a given power, boxcar shaped transmissions are ideal [203]. Finding systems with these properties is unfortunately not easy. We discuss below the different strategies proposed so far.

### v2

# Quantum thermocouples: nonlocal conversion and control of heat in nanostructures

@@ I Introduction
The thermocouple is the minimal thermoelectric device able to use the Seebeck effect [1] to transduce a heat flow into an electric response measurable in an isothermal conductor. It is useful for the autonomous conversion of environmental (or waste) heat into useful power [2, 3], or as a temperature sensor when coupled to a voltmeter. The basic principle of a macroscopic thermocouple is sketched in Fig. 1(a). Two junctions separate the conductor, at a base temperature $T_{C}$, from a region that is in contact with a heat source at temperature $T_{H}$. If the two junctions have opposite thermoelectric response (e.g., one being n-type and the other p-type), the electron-hole excitations, thermalized at a temperature $T_{M}$ in the central region, are spatially separated to opposite contacts. This can then either generate a current or be measured as a voltage difference which depends on the difference of the linear Seebeck coefficient of each junction: $S_{p}-S_{n}$. The reverse process is also possible: driving a current through the conductor can be used to extract heat from the central region, in which case one has a Peltier refrigerator.
@@
Reducing the size of the thermocouple to nanometer scales [4] comes with a number of key differences. The first to be noted was that the reduced dimensionality provides the required spectral properties for efficient thermoelectric response [5, 6], which can also be easily tuned with gate voltages, as illustrated in Fig. 1(b) for a resonant-tunneling barrier. At this scale, it becomes difficult to maintain large temperature differences. However, one benefits from the system response easily becoming nonlinear. From the point of view of the physical processes, the size of the region in contact with the heat source can be of the order of or smaller than the carrier thermalization length. This has two possible consequences: the thermoelectric response may become nonlocal [7, 8] (not determined by the series response of the junctions) or determined by nonthermal states [9], as represented in Fig. 1(c). Finally, at low enough temperatures quantum properties arising from electrons propagating as waves in conductors smaller than the phase coherence length start to contribute, leading to interference effects [10, 11]. Also, single particle interactions (electron-electron or electron-photon) can be relevant [12]. This is the regime where we focus our interest in this review. Properties from quantum materials (e.g., superconductors or edge states in quantum Hall regimes, spin-orbit couplings) can also be exploited [8, 13, 14, 15].
@@
While the electronic transport through nanostructures has been intensively investigated during the last decades both experimental [16] and theoretically [17], progress in the control of thermal and thermoelectric flows was somehow delayed, mainly due to the difficulty in controlling and measuring minute temperature differences in an experiment and in eliminating phonon drag contributions. This has however changed in the last years [18, 19, 20, 21]. The interest is not restricted to the possibility to generate useful thermodynamic outputs but also to use the thermoelectric response as sensitive probes for transport mechanisms, in particular involving the difficult to measure energy flow. Controlled thermoelectric measurements have been achieved in 2DEGs [22, 23, 24, 25], quantum point contacts [26, 27, 28, 29], quantum dots [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], single-electron transistors [43], nanowires [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], scanning probes [56, 57], molecular junctions [58, 59, 60, 61, 62], superconducting tunnel junctions [63, 64, 65] or magnetic wires [66], also in cold atom gases [67]. In particular, three-terminal quantum dot arrangements in thermocouple geometries have been realized using electron-electron interactions [68, 69, 70], hot cavities [71], electron-phonon [72, 73] or electron-photon interactions [74, 75] to couple the conductor to the heat source. Multiterminal quantum Hall edge states [76, 77, 78], (local) Joule heating in graphene [79] and normal-superconductor interfaces [80] have also been considered. Interestingly, similar responses have been measured when coupling the conductor to nonequilibrium fluctuations [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], which can be related to energy quanta emitted by relaxing electrons [95].
@@
This progress has been accompanied by theoretical works proposing the multiterminal thermoelectric response to take advantage of the separation of thermal and particle currents [96, 97, 98] or to access properties not available in the electrical conductance, for instance nonabelian statistics of carriers [99], odd-frequency superconducting correlations [100], crossed-Andreev reflection [101, 102, 103], correlations between edge states [104, 105] or details of interfaces between superconductors and topological materials [106, 107, 108]. At the same time, questions related to quantum thermodynamics [109], e.g. the efficiency of quantum heat engines [110, 111], or the role of fluctuations [112, 113, 114, 115, 116, 117] and of information [96, 118, 119, 120, 121, 122, 123, 124] have found in the thermoelectric effect a suitable platform to be addressed.
@@
Particularly relevant to our discussion in this review is the configuration where a quantum system is coupled to two electronic terminals at the same temperature, with a third terminal (or bath) injecting only heat. Under the appropriate broken symmetries, the injected heat is rectified into a particle current. Initial proposals considered electron-phonon coupling [125] and electron-electron interactions [96] as possible mechanisms for the transfer of heat (later realized experimentally [68, 69, 70, 72]), followed by configurations exploiting inelastic scattering [126] (realized with quantum dots [71]), magnons [127] or photons [128, 129, 130, 131]. The variety of transport configurations considered is large, including capacitively coupled quantum dots [96, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141], molecules [125, 142, 143, 144], nanowires [145], resonant tunneling [126, 146, 147, 148, 149, 150, 151, 152], edge states [153, 154, 155, 156, 157, 158, 159, 160, 104, 161], beam splitters [162, 163, 115, 164, 165], quantum dot arrays [166, 167, 168, 169, 170, 97, 171, 172, 173, 174, 175], and tunnel [128, 176, 177], p-n [178] or Josephson [179] junctions. Three terminal geometries have also been considered for exploring the thermoelectric efficiency bounds when one of the terminals is a probe [180, 181, 182, 183, 184, 185, 186, 187]. In the case that the resource consists on several terminals, nonthermal configurations [135, 159] or thermal drag effects [188, 189, 190, 191] can be exploited.
@@ II Multiterminal particle and heat transport
We are interested in configurations as the one depicted in Fig. 1(c), with multiple (more than two) terminals allowing for particle ($I_{i}$) and heat ($J_{i}$) currents, which will be positive when flowing out of their reservoirs, labeled by $i$ or $j$. The system is taken out of equilibrium by differences of the reservoir electrochemical potentials, $\mu_{i}$, and temperatures, $T_{i}$. Allowing the system to be composed of electrically isolated regions (i.e., in the case of having capacitively-coupled conductors), labeled by $\alpha$, particle and energy conservation are expressed as $\sum_{i\in\alpha}I_{i}=0$ and $\sum_{i}J_{i}+\sum_{\alpha}P_{\alpha}=0$, where $P_{\alpha}=-\sum_{i\in\alpha}\mu_{i}I_{i}$ is the dissipated power in conductor $\alpha$. The system will perform as a thermodynamic engine when power is generated ($P_{\alpha}>0$) in some conductor $\alpha$, as a refrigerator when heat flows out of a cold reservoir ($J_{i}>0$), or as a heat pump when heat flows into a hot reservoir ($J_{i}<0$). In a multiterminal setup, whether a terminal is considered “hot” or “cold” depends on its temperature being larger or smaller than a certain reference temperature $T_{0}$ [193]. Differently from macroscopic thermocouples though, fluctuating systems that are smaller than the electronic thermalization length may contain regions without a well defined temperature [135].
Our discussion is not restricted to the linear response regime, however considering this regime gives useful intuition and will be invoked here when appropriate. Linearizing the particle and heat currents with respect to variations of the electrochemical potentials and temperatures, we get the elements of the Onsager matrix:
@@
Multiterminal heat engines have obviously multiple possible resources and multiple possible thermodynamically useful outputs, that can even occur simultaneously [147, 198, 199, 159, 200, 193, 160, 201, 174, 202, 174, 203]. Hence, the usual definition of efficiency in terms of the ratio of output power and absorbed heat current needs to be revisited. Think for instance in the configuration of Fig. 1(c): uncareful definitions of the efficiency may diverge when power is generated by a current between terminals $L$ and $R$, while the heat absorbed by the system, $J_{H}+J_{C}$, vanishes [159]. To have a thermodynamically consistent definition of the engine performance, it is useful to consider the change in the nonequilibrium free energy of the reservoirs, $\dot{F}_{i}=-I_{i}\mu_{i}+J_{i}(T_{0}-T_{i})/T_{i}$ [193, 160], with the second law imposing that $\sum_{i}\dot{F}_{i}\leq 0$, i.e., free energy is consumed in dissipation. The efficiency is then written by selecting the positive electric and thermal contributions and dividing them by the negative ones [193]:
@@
where $\sum_{i}^{\pm}x_{i}=\sum(x_{i}\pm|x_{i}|)/2$. Generated power is positive, dissipated power is consumed therefore negative. Extracting/injecting heat from/into cold/hot reservoirs (i.e., cooling/pumping) are positive, damped heat is negative. [^1] Whether a reservoir is considered hot or cold is is established by the reference temperature, $T_{0}$, which hence has an operational meaning. As an example, consider the basic three-terminal thermocouple, with one hot reservoir injecting heat but no particles ($I_{H}=0$, $J_{H}>0$) and the other two terminals being at the same temperature ($T_{L}=T_{R}=T_{0}$) and supporting a particle current ($I_{L}+I_{R}=0$). Then if power is generated by the current flowing against the electrochemical potential difference, $I_{L}(\mu_{R}-\mu_{L})$, we recover $\eta=P/(J_{H}\eta_{C})$, with the Carnot efficiency $\eta_{C}=1-T_{0}/T_{H}$. Other definitions of the efficiency are possible based on the rate of entropy change of the different reservoirs [174] that only involve temperatures of the baths.
@@ III Noninteracting electrons
A simple and elegant description of multiterminal quantum transport is given by scattering theory [205], as long as interactions admit a mean field approach [206], see Refs. [207, 208, 8, 10, 11] for more details. Reservoirs at different electrochemical potentials $\mu_{i}$ and temperature $T_{i}$ inject electrons into the terminals of the conductor, which is treated as a wave scatterer. The scattering matrix ${\cal S}_{ji}(E)$ represents the amplitude for electrons injected with energy $E$ from terminal $i$ to be transmitted and absorbed by reservoir $j$. The initial success of the theory originated in the description of quantized conductance [209, 210] and the understanding of the properties and symmetries of the linear response regime [211, 212, 213]. However, particle ($I_{i}$) and heat ($J_{i}$) currents can be more generally defined in terms of the occupation of the different terminals, given by the fermionic reservoir distribution function, $f_{i}(E)=1/\{1+\exp[(E-\mu_{i})/k_{\text{B}}T_{i}]\}$:
$I_{i}$ $=2h^{-1}\sum_{j}\int{dE}{\cal T}_{ij}(E)[f_{i}(E)-f_{j}(E)]$ (9)
$J_{i}$ $=2h^{-1}\sum_{j}\int{dE}(E-\mu_{i}){\cal T}_{ij}(E)[f_{i}(E)-f_{j}(E)],$ (10)
@@
where $h$ is the Planck constant, the factor 2 is due to spin degeneracy and ${\cal T}_{ij}(E)=|{\cal S}_{ij}(E)|^{2}$ is the transmission probability for an electron injected from terminal $j$ and absorbed by terminal $i$. The unitarity of the scattering matrix preserves particle and energy conservation [8] as expressed in Sec. II. With this convention (particle and heat currents are positive when flowing out of reservoirs) and linearizing Eqs. (9) and (10) when $(\mu_{i}-\mu)/k_{\text{B}}T$ and $(T_{i}-T)/T\ll 1$, the linear coefficients in Eq. (7) can be written as $G_{ij}=g_{ij}^{(0)}$, $L_{ij}=g_{ij}^{(1)}/T$, $M_{ij}=g_{ij}^{(1)}$, $K_{ij}=g_{ij}^{(2)}/T$, with [214]
$g_{ij}^{n}=\frac{2}{h}\int{dE}(E-\mu)^{n}[N_{i}\delta_{ij}-{\cal T}_{ij}(E)]\left(-\frac{\partial f_{0}(E)}{\partial E}\right),$ (11)
@@
where $N_{i}$ is the number of channels of the lead connecting the system to terminal $i$ and $f_{0}(E)$ is the Fermi distribution at equilibrium. In the limit $\mu\gg k_{\text{B}}T$, these expressions show that the thermoelectric response requires that the scattering problem includes some explicit energy dependence i.e., that the nanostructure acts as a filter to break electron-hole symmetry. If the transmissions are energy-independent, ${\cal T}_{ij}(E)=\tau_{ij}$, one gets $G_{ij}=2(N_{i}\delta_{ij}-\tau_{ij})/h$ and $K_{ij}=2q_{H}(N_{i}\delta_{ij}-\tau_{ij})$, with the quantum of thermal conductance, $q_{H}=\pi^{2}k_{\text{B}}^{2}T/3h$ [215], for the conductances, and no thermoelectric response: $L_{ij}=M_{ij}=0$. The fulfillment of the Onsager reciprocity relations is hence related to the symmetries of the scattering matrix under time reversal satisfying microreversibility: ${\cal S}_{ij}(E,B)={\cal S}_{ji}(E,-B)$ [216], and therefore also ${\cal T}_{ij}(B)={\cal T}_{ji}(-B)$ [213]. The linear coefficients plotted in Fig. 1(b) correspond to a Lorentzian transmission probability usually found in quantum dots [30, 31, 35]. Nonlinear contributions require a careful treatment of the local potentials [217, 218], determined by particle and entropic injectivities and the onset of interaction effects [219, 220, 221, 222, 223, 224, 225, 226].
Let us illustrate the thermoelectric effect with the most relevant case for the remaining discussion: a three-terminal thermocouple, with terminals $L$ and $R$ forming the conductor where current is to be generated, and terminal $H$ on average injecting only heat. The later is described by imposing $H$ to be a voltage probe [227, 228]: i.e., the chemical potential $\mu_{H}$ will adapt to satisfy the boundary condition $I_{H}=0$, while having its temperature fixed, $T_{H}$. We can then separate the current in two terms, $I_{i}=I_{i,el}+I_{i,inel}$, depending on whether electrons flow elastically between the conductor terminals $I_{i,el}\propto\int{dE}{\cal T}_{LR}(E)[f_{L}(E)-f_{R}(E)]$, or experience inelastic scattering at $H$: $I_{i,inel}=(2/h)\int{dE}{\cal T}_{iH}(E)[f_{i}(E)-f_{H}(E)]$. Assume for simplicity that the conductor is in equilibrium ($\mu_{L}=\mu_{R}$ and $T_{L}=T_{R}$), so we write $f_{L}(E)=f_{R}(E)=f_{0}(E)$. Then, transport is fully characterized by the difference of occupation of the heat source and the conductor, $f_{H}(E)-f_{0}(E)$, and the scattering probabilities ${\cal T}_{ij}(E)$. In that case, using $\sum_{i}{\cal T}_{ij}=\sum_{j}{\cal T}_{ij}=1$ and particle conservation, $I\equiv I_{L}=-I_{R}$, we are left with only two equations:
$\begin{aligned} I&=\frac{1}{h}\int dE[\mathcal{T}_{LH}(E)-\mathcal{T}_{RH}(E)][f_{0}(E)-f_{H}(E)]\\ 0&=\int dE[\mathcal{T}_{HL}(E)+\mathcal{T}_{HR}(E)][f_{H}(E)-f_{0}(E)],\end{aligned}$ (12)
@@
corresponding to the generated current and the probe condition, respectively. Note that the elastic transmission, ${\cal T}_{LR}$, does not contribute, and the generated zero bias current is entirely due to the inelastic scattering of electrons at the probe terminal. By inspection of Eq. (12), we see that for the system to work as a thermocouple able to generate a finite current, $I\neq 0$, it is required that the two transmissions are energy-dependent (in order to generate a thermoelectric response via the broken electron-hole symmetry), and that this dependence also breaks inversion (left-right) symmetry, i.e., $\mathcal{T}_{LH}(E)\neq c\mathcal{T}_{RH}(E)$, with $c$ a positive number. Otherwise, the two integrals become proportional, so the probe condition imposes $I=0$. This implies that the symmetry breaking must be a property of the conductor, not of the coupling to the heat source.
@@
To generate a finite power, the particle current must flow against a finite bias, $\mu_{L}\neq\mu_{R}$. In this case, the elastic transmissions ${\cal T}_{LR}(E)$ and ${\cal T}_{RL}(E)$ introduce contributions that are independent of the heat source (assuming that we can neglect nonlinear effects on the internal potential of the conductor [229]) and hence are detrimental to the device performance [192]. Therefore power will be enhanced when the heat source is connected in series between the two conductor terminals [126], so all transported electrons are forced to thermalize at the probe. In analogy with the macroscopic thermocouple, the effect is enhanced if the two contributions to the current in Eq. (12) have opposite signs, i.e., one is dominated by electrons and the other by holes. This can be done by placing scatterers that separate the heat source from the two conductor terminals having sufficiently sharp spectral features antisymmetric with respect to the electrochemical potential, e.g., quantum dots [126]. It was found that in order to optimize the efficiency for a given power, boxcar shaped transmissions are ideal [230], also for nonthermal resources [231]. Finding systems with these properties is unfortunately not easy, so alternatives are requested [232]. We discuss below the different strategies proposed so far. To ease the discussion, we will restrict to single-channel and time reversal symmetric conductors in the following, except when explicitly including a magnetic field.